OpenStudy (anonymous):

calculate the pH of a solution prepared by dissolving 0.3 moles of a weak organic acid (HA) and 0.40 moles of sodium salt of this acid (NaA) in 1.00 L of solution (ka of HA = 7.5 * 10 ^-5)

OpenStudy (rogue):

There are both the weak acid and its conjugate base in the solution, so it's a buffer. The pH of a buffer can be found using the Hender-Hasselbalch equation.\[pH = pK_a + \log_{10} ( \frac {[A^-]}{[HA]})\]Because the A- and HA are in the same 1 L solution, the mole ratio and concentration ratio is the same, so we don't have to calculate for their concentrations.\[pH = - \log_{10} (7.5 \times 10^{-5}) + \log_{10} ( \frac {0.40 \space mol \space A^-}{0.30 \space mol \space HA}) = 4.25\]